Asteelcompanyisproducingsteelshaftsforaxisofturbines by  forgingprocess. The forgingtemperatureis1200  °C . After theforgingprocesstheshafthastobecooleddown . This isdonebysprayingwaterhomogeniouslyontothesurfaceoftheshaft.

I.-Describethecoolingprocesswiththehelpof dimensionlessnumbers. Makethefollowingassumptions:

- The shaftisa cylinderwithdiamterD andlenghtl

- The physical propertiesofthe steelare:

    - density ofthesteel:  ρ

    - thermal diffusity: a

-homogenoustemperatureat thebeginning: T_0,

-constantsurfacetemperatureduringcooling: T_1

-meantemperatureinside: T(t)

II.- How will the cooling time change  when the mass of the shaft will be halved (shape of the shaft will be unchanged: L/D=const)?
Answer #1

Diameter of shaft (d) = D. Length of shaft (2) = 1 Forging temperature (fo) = 1200°C - Density of steel = p. Thermal diffusivbe no Variation position. In this of Temperature with case we can. write, -has (T-To) = 4 me (IT) Solving we get (T-Tool - e-D Let Initial mass = m . Final mass = m/ (b) = constant = c (say) - 0 we know, we mass = (Density & Volume) If we are not cha2 + c) from equation (1) t, (% + c)² ta (1/2 + c)2 - se, Also from equation ① and on lix Cils) = 2* (1/)24 17 - l = 2 & 3 Fro

All answers are boxed and steps are shown clearly. Thank You and have a great day.